Let X-(Xt ,0 < t < 1} be an arithmetic Brownian motion starting from 0 with drift parameter μ-0.2 and variance parameter ơ2-0.125. 1. Calculate the probability that X2 is between 0.1 and 0.5 2. Given that X 0.6, find the probability that X2 is between 0.1 and 0.5 3. Given that Xi- 0.2, find the covariance between X2 and X3
Answer #1

un thatXtis Bow ow nion motion 0-2 t +-0.35-B(t) = 0-125:t Vo 3 -PC05 5く乙く0-1851 ニ0-5734 -0:2911 = 0.2823Know thatfox bownian mohor v1x2/X3 ะ 0.6 )-(3-2) 2-0 : 661 3 No. 667 Vo 667 P (-0.367/ZZC 0122) s o 518 SS-.3 5677 는 0:1917 To Fir Cov Co 2 t t0 35 3C2), 0.2 t +0.35B(3) o 35 (ou (B(22,83)and also Ionditional disbuibution DF ou nian mothon is same d unienditioal dis Hhihuthon f brownian mata in се